ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 226, 2018

Alpay, D., M. E. Luna-Elizarrarás, M. Shapiro and D. Struppa Gleason’s problem, rational functions and spaces of left-regular functions: The split-quaternion setting

319

Angeleri Hügel, L., J. Šaroch and J. Trlifaj Approximations and Mittag-Leffler conditions the applications

757

Benoist, Y. and J.-F. Quint On the regularity of stationary measures

1

Bilu, Y. and J. Gillibert Chevalley–Weil theorem and subgroups of class groups

927

Bochi, J., Ch. Bonatti and K. Gelfert Dominated Pesin theory: convex sum of hyperbolic measures

387

Bonatti, Ch. See J. Bochi De Commer, K., P. Kasprzak, A. Skalski and P. M. Sołtan

Quantum actions on discrete quantum spaces and a generalization of Clifford’s theory of representations

475

Conde-Alonso, J. M., J. Parcet and É. Ricard On spectral gaps of Markov maps

189

Coskun, I. and J. Huizenga The nef cone of the moduli space of sheaves and strong Bogomolov inequalities

205

Cummings, J., S.-D. Friedman, M. Magidor, A. Rinot and D. Sinapova Ordinal definable subsets of singular cardinals

781

Detomi, E., M. Morigi and P. Shumyatsky On profinite groups with word values covered by nilpotent subgroups

993

Donoso, S. and W. Sun Quantitative multiple recurrence for two and three transformations

71

Ellis, P. The classification problem for simple unital finite rank dimension groups

419

Feng, D.-J. and Y. Xiong Affine embeddings of Cantor sets and dimension of αβ-sets

805

Filip, S. Families of K3 surfaces and Lyapunov exponents  29

Fraser, J. M. Distance sets, orthogonal projections and passing to weak tangents

851

Friedman, S.-D. See J. Cummings

Gehrmann, L. On Shalika models and p-adic L-functions  237

Gelfert, K. See J. Bochi Giambruno, A. and M. Zaicev Central polynomials and growth

functions

15

Gillibert, J. See Y. Bilu Ginzburg, D. Non-generic unramified representations in metaplectic

covering groups

447

Goldbring, I. and M. Lupini Model-theoretic aspects of the Gurarij operator system

87

Groves, D. and H. Wilton The structure of limit groups over hyperbolic groups

119

Hàn, H. and A. Jiménez Maximum number of sum-free colorings in finite abelian groups

505

Huizenga, J. See I. Coskun Ito, T. and T. Tsushima Cuspidal representations in the cohomol-

ogy of Deligne–Lusztig varieties for GL(2) over finite rings

877

Ivanov, S. O. and R. Mikhailov On lengths of Hℤ-localization towers

635

Jiménez, A. See H. Hàn

Joseph, A. Convexity properties of the canonical S-graphs  827
Kanigowski, A. Slow entropy for some smooth flows on surfaces  535

Kanigowski, A. and A. V. Solomko On isomorphism problem for von Neumann flows with one discontinuity

685

Kasprzak, P. See K. De Commer

Lechner, R. Factorization in SL^{∞}  957

Len, Y. and D. Ranganathan Enumerative geometry of elliptic curves on toric surfaces

351

Luna-Elizarrarás, M. E. See D. Alpay Lupini, M. See I. Goldbring Magidor, M. See J. Cummings Mikhailov, R. See S. O. Ivanov Morigi, M. See E. Detomi Pacelli Bessa, G., L. F. Pessoa and M. Rigoli Vanishing theo-

rems, higher order mean curvatures and index estimates for selfshrinkers

703

Parcet, J. See J. M. Conde-Alonso Pessoa, L. F. See G. Pacelli Bessa

Pošta, P. Dirichlet problem and subclasses of Baire-one functions  177

Quint, J.-F. See Y. Benoist Ranganathan, D. See Y. Len Ricard É. See J. M. Conde-Alonso Rigoli, M. See G. Pacelli Bessa Rinot, A. See J. Cummings

Šaroch, J. Approximations and Mittag-Leffler conditions The tools  737

Šaroch, J. See L. Angeleri Hügel Shapiro, M. See D. Alpay Shlapentokh, A. First-order decidability and definability of integers

in infinite algebraic extensions of the rational numbers

579

Shumyatsky, P. See E. Detomi Sinapova, D. See J. Cummings Skalski, A. See K. De Commer Solomko, A. V. See A. Kanigowski Sołtan, P. M. See K. De Commer Struppa, D. See D. Alpay Sun, W. See S. Donoso Trlifaj, J. See L. Angeleri Hügel Tsushima, T. See T. Ito Wang, P. and J. Zhuang Convexity of level lines of maximal space-

like hypersurfaces in Minkowski space

295

Wilton, H. See D. Groves Xiong, Y. See D.-J. Feng Zaicev, M. See A. Giambruno Zhuang, J. See P. Wang